# Generation of large-scale magnetic fields upstream of gamma-ray burst afterglow shocks

### v1

@@ Abstract
The origins of the magnetic fields that power gamma-ray burst (GRB) afterglow emission are not fully understood. One possible channel for generating these fields involves the pre-conditioning of the circumburst medium: in the early afterglow phase, prompt photons streaming ahead of the GRB external shock can pair produce, seeding the upstream with drifting electron-positron pairs and triggering electromagnetic microinstabilities. To study this process, we employ 2D periodic particle-in-cell simulations in which a cold electron-proton plasma is gradually enriched with warm electron-positron pairs injected at mildly relativistic speeds. We find that continuous pair injection drives the growth of large-scale magnetic fields via filamentation-like instabilities; the temporal evolution of the field is self-similar and depends on a single parameter, $\left[\alpha/(t_{f}\omega_{pi})\right]^{1/2}t\omega_{pi}$, where $\alpha$ is the ratio of final pair beam density to background plasma density, $t_{f}$ is the duration of pair injection, and $\omega_{pi}$ is the plasma frequency of background protons. Extrapolating our results to parameter regimes realistic for long GRBs, we find that upstream pair enrichment generates weak magnetic fields on scales much larger than the proton skin depth; for bright bursts, the extrapolated coherence scale at a shock radius of $R\sim 10^{17}\,{\rm cm}$ is $\lambda_{y}\right>\sim 100\;c/\omega_{pi}$ and the corresponding magnetization is $\sigma\sim 10^{-8}$ for typical circumburst parameters. These results may help explain the persistence of magnetic fields at large distances behind GRB shocks.
@@ 1 Introduction
Our uncertainty regarding the nature of GRB afterglows is intimately tied to our incomplete understanding of relativistic, weakly magnetized, collisionless shocks (e.g., Sironi et al. 2013, 2015; Vanthieghem et al. 2020). The low densities of the environments surrounding GRB progenitors yield Coulomb mean free paths that far exceed the system size, meaning that GRB external shocks are mediated not by binary particle collisions but by collective plasma interactions (Moiseev & Sagdeev, 1963; Medvedev & Loeb, 1999; Gruzinov & Waxman, 1999; Lyubarsky & Eichler, 2006; Brainerd, 2000; Wiersma & Achterberg, 2004; Kato, 2007). The ambient magnetic fields of circumburst environments are typically very weak, with $\sigma_{0}\equiv B_{0}^{2}/4\pi n_{i}m_{i}c^{2}\sim 10^{-9}-10^{-5}$, where $n_{i}$ is the proton density, $m_{i}$ is the proton mass and $c$ is the speed of light. In this weakly magnetized, collisionless regime, we expect shocks to be mediated by magnetic turbulence produced by the filamentation (or Weibel) instability (Weibel, 1959; Fried, 1959), which channels the free energy of an anisotropic flow into magnetic energy, effectively generating magnetic fields from scratch (Achterberg & Wiersma, 2007; Achterberg et al., 2007; Bret et al., 2010b; Lemoine & Pelletier, 2010, 2011; Rabinak et al., 2011; Nakar et al., 2011; Shaisultanov et al., 2012). Unfortunately, the evolution of this magnetic turbulence is highly nonlinear – with the magnetic fields backreacting on the same particles that produce them – making Weibel-mediated shocks nearly intractable to study analytically (Kirk et al., 2000; Keshet et al., 2009a, b; Bret et al., 2013, 2014; Pelletier et al., 2019; Lemoine et al., 2019c, b, a); only fully kinetic numerical simulations can completely capture these nonlinear interactions (Spitkovsky, 2005; Birdsall & Langdon, 1991; Matsumoto, 1993), but even these simulations are strained by limited computational resources, leaving many questions still unanswered (Keshet et al., 2009a).
@@
A crucial question in the modeling of GRB afterglows concerns the extent of the afterglow emission region. If the magnetic fields powering the synchrotron afterglow are indeed generated by the filamentation instability, then theory predicts that these fields should grow on scales comparable to the proton skin depth (Medvedev & Loeb, 1999), leading to fast decay downstream of the shock (Chang et al., 2008; Gruzinov, 2001; Lemoine, 2015); instead, observations combined with analytical models suggest that these downstream fields may extend up to $10^{8}$ proton skin depths behind the shock (Waxman, 2006). In addition, if particles at the shock were to scatter in exclusively microscale fields, their non-thermal acceleration would be limited (Sironi et al., 2013; Reville & Bell, 2014; Huang et al., 2022). One possible solution to this problem involves the feedback of particles accelerated at the shock (Keshet et al., 2009a; Grošelj et al., 2024): as particles are accelerated to higher and higher energies via the Fermi process (Blandford & Eichler, 1987), they will penetrate farther into the upstream, seeding progressively larger-scale magnetic fluctuations that decay more slowly and scatter particles more efficiently. An alternative solution – particularly relevant to long bursts, in which the GRB’s external shock propagates through the dense wind of the Wolf-Rayet progenitor star – involves the loading of the circumburst medium with electron-positron pairs prior to the passage of the shock (Thompson & Madau, 2000; Beloborodov, 2002; Kumar & Panaitescu, 2004; Ramirez-Ruiz et al., 2007; Derishev & Piran, 2016; Grošelj et al., 2022). Early in the afterglow phase (i.e., within a few minutes of the burst), prompt photons stream just ahead of the shock front, where a small fraction could scatter off of electrons in the upstream plasma; the scattered photons are then susceptible to collisions with other outgoing prompt photons, producing warm electron-positron pairs. [^1] The photons required for upstream pair production may alternatively be sourced by synchrotron radiation downstream of the shock (Derishev & Piran, 2016). As each pair increases the optical depth of the upstream medium – thus increasing the probability of photon scattering – the rate of pair creation increases exponentially, ultimately resulting in a pair density profile that decreases away from the shock front (Beloborodov et al., 2014). The streaming of these electron-positron pairs through the circumburst plasma should trigger the filamentation instability well before the shock arrives, potentially pre-seeding the upstream with magnetic fields; if these fields are coherent on sufficiently large scales, they could survive far into the downstream.
@@
In this paper, we overcome the limitations of Garasev & Derishev (2016) by carrying out periodic, 2D, fully kinetic simulations of the continuous enrichment of an electron-proton plasma (with realistic mass ratio) by relativistic electron-positron pairs carrying net momentum in the fluid frame. We also explore the physics of this system in depth, identifying trends that allow us to extrapolate our simulation results to regimes realistic for long GRBs. We conduct an extensive parameter scan over two key quantities: $t_{f}$ – the pair injection duration (or, equivalently, the time for the upstream fluid element to encounter the shock front) – and $\alpha$ – the ratio of the final pair beam density to the background plasma density. In the range of $t_{f}$ and $\alpha$ that we explore, we find that continuous pair injection drives the growth of large-scale magnetic fields via filamentation-like instabilities; the temporal evolution of the system is self-similar and depends on a single parameter, $\left[\alpha/(t_{f}\omega_{pi})\right]^{1/2}t\omega_{pi}$, where $\omega_{pi}$ is the plasma frequency of the background protons. At the end of the pair injection phase, the mean wavenumber of the magnetic field (in units of $\left(c/\omega_{pi}\right)^{-1}$) scales approximately with $\alpha^{-1/12}\,\left(t_{f}\omega_{pi}\right)^{-2/3}$ and the self-generated magnetic energy density (in units of the proton rest mass energy density) scales with $\alpha^{5/8}\,\left(t_{f}\omega_{pi}\right)^{-3/4}$. Using these trends to extrapolate to realistic regimes, we find that upstream pair enrichment generates weak magnetic fields on scales much larger than the proton skin depth; for bright bursts, the extrapolated coherence scale at a shock radius of $R\sim 10^{17}\,{\rm cm}$ is $\lambda_{y}\right>\sim 100\;c/\omega_{pi}$ and the corresponding magnetization is $\sigma\sim 10^{-8}$ for typical circumburst parameters. Since larger-scale fields decay slower, our results may help explain the persistence of magnetic fields at large distances behind GRB shocks.
This paper is structured as follows: in Section 2, we detail our simulation setup; in Section 3, we analyze an example simulation, highlighting the salient features of the magnetic field evolution; in Section 4, we compare how the magnetic field evolution varies with $\alpha$ and $t_{f}$; in Section 5, we extrapolate our simulation results to realistic regimes; and in Section 6, we summarize our main findings and conclude. We provide supplementary information in the appendices: in Appendix A, we compare our 2D results with a 3D simulation; in Appendix B, we compare simulations with varying beam Lorentz factors; and in Appendix C, we show that the results presented in this paper are converged numerically.
@@ 2 Simulation Setup
We employ the electromagnetic particle-in-cell (PIC) code TRISTAN-MP (Spitkovsky et al., 2019; Spitkovsky, 2005), which is especially well-suited for handling relativistic flows in collisionless plasmas. The code solves the Maxwell-Vlasov system using a second-order Finite-Difference Time-Domain (FDTD) scheme on a standard Yee mesh (Yee, 1966) and pushes particles (with first-order shape functions) using the Boris algorithm (P., 1970). The electromagnetic fields are extrapolated to particle positions using a bilinear interpolation function (or a trilinear interpolation function for our 3D run). At each time step, after depositing the electric current to the grid using the charge-conserving zig-zag scheme (Umeda et al., 2003), we apply twenty passes of a 3-point (1-2-1) digital current filter in each direction to smooth out non-physical short-wavelength oscillations.
@@
We run fully periodic 2D simulations (in the $xy$-plane) of a cold, initially unmagnetized electron-proton plasma (the “background”) that is gradually enriched with warm, mildly relativistic electron-positron pairs (the “beam”). The background particles and the beam particles are both distributed homogeneously throughout the simulation box. The background is initially at rest and follows a Maxwellian with $\frac{k_{B}T_{i}}{m_{i}c^{2}}=\frac{m_{e}}{m_{i}}\frac{k_{B}T_{e,\,bg}}{m_{e}c^{2}}=10^{-7}$ (where $k_{B}$ is Boltzmann’s constant, $c$ is the speed of light, $T_{i}$ is the proton temperature, $T_{e,\,bg}$ is the background electron temperature, $m_{i}$ is the proton mass, $m_{e}$ is the electron mass, and $m_{i}/m_{e}=1836$). The beam pairs are drawn from a Maxwell-Jüttner distribution with $\frac{k_{B}T_{b}}{m_{e}c^{2}}=1$ (for $T_{b}$ the initial temperature of both beam species), drifting along the positive $x$ direction with bulk Lorentz factor $\gamma_{b}=1.5$.
@@
The beam particles are injected in pairs at a constant rate throughout the duration of the simulation. The injection rate is determined by two key parameters: $t_{f}$ – the injection duration – and $\alpha$ – the ratio of the total beam density to the total background density at the end of injection. In terms of these quantities, $n_{bg}\times\frac{\alpha}{t_{f}\omega_{pb}}$ electron-positron pairs are injected per plasma time ($\omega_{pb}^{-1}$). Here, $n_{bg}$ is the apparent number density of the background plasma (i.e., the total number density of the background electrons and the protons) and $\omega_{pb}$, the final beam electron plasma frequency, is defined to be
$\omega_{pb}\equiv\sqrt{\frac{4\pi(n_{b}/2)e^{2}}{m_{e}}},$ (1)
@@
for $n_{b}$ the total beam density at the end of pair injection. It is also useful to define the plasma frequency of the background protons,
$\omega_{pi}\equiv\sqrt{\frac{4\pi(n_{bg}/2)e^{2}}{m_{i}}}=\omega_{pb}\sqrt{\frac{m_{e}}{m_{i}\alpha}}.$ (2)
@@
In our simulations, all quantities are normalized to the properties of the pair beam at the end of injection to ensure that all length-scales are well-resolved by our grid; since we only consider $\alpha\geq 1$, the skin depth of the protons ($c/\omega_{pi}$) and the skin depth of the background electrons ($c/\omega_{pe}\equiv\sqrt{m_{e}/m_{i}}\,c/\omega_{pi}$) are always at least as large as the skin depth of the beam electrons. Throughout this paper, we cite results both in units of the beam electrons and in units of the background protons.
@@
In our suite of simulations, we consider values of $\alpha$ equal to 1, 2, 4, 8, and 16; for each $\alpha$, we vary $t_{f}$ among $t_{f}\omega_{pb}=5\times 10^{3}$, $10^{4}$, $2\times 10^{4}$, and $4\times 10^{4}$ (respectively corresponding to $t_{f}\omega_{pi}\sim 117/\sqrt{\alpha}$, $233/\sqrt{\alpha}$, $467/\sqrt{\alpha}$, and $934/\sqrt{\alpha}$). These values are small compared to those appropriate for real GRBs, since our realistic mass ratio prevents us from directly simulating larger $\alpha$ and $t_{f}$. However, our range of $\alpha$ and $t_{f}$ is broad enough to allow for reliable extrapolation to higher values.
@@
For our production runs, we primarily use 36 final beam particles per cell per species; the number of background particles per cell per species is thus $36/\alpha$. For a few representative cases, we also use 72 final beam particles per cell per species (and $72/\alpha$ background particles per cell per species) to check for numerical convergence (see Appendix C). We resolve the final beam electron skin depth, $c/\omega_{pb}$, using 10 cells. We employ square simulation boxes with side lengths of either 600 $c/\omega_{pb}$ (6000 cells) or 1000 $c/\omega_{pb}$ ($10^{4}$ cells); we use the larger boxes for our $t_{f}\omega_{pb}=2\times 10^{4}$ and $4\times 10^{4}$ cases to ensure that the dominant scale of the magnetic field is not artificially limited by the box size. For our $\alpha=16$ simulations (i.e., the cases with the largest proton skin depth), $c/\omega_{pi}\sim 200\;c/\omega_{pb}$, so our simulation boxes always fit at least three proton skin depths in each direction.
@@
To check our 2D results, we run a single 3D simulation with our fiducial parameters, $\alpha=2$ and $t_{f}\omega_{pb}=10^{4}$; this simulation is analyzed in Appendix A. For this 3D run, we use 18 final beam particles per cell per species, we resolve $c/\omega_{pb}$ with 10 cells, and we employ a cubic simulation box with a 100 $c/\omega_{pb}$ (1000 cell) side length.
@@
In Appendix C, we show that our simulations converge numerically with respect to both particles per cell and spatial resolution, and we demonstrate that the growth of the magnetic field is not impeded by the size of our simulation box.
3 Case study: $\alpha=2,\;t_{f}\omega_{pb}=10^{4}$

### v2

@@ Abstract
The origins of the magnetic fields that power gamma-ray burst (GRB) afterglow emission are not fully understood. One possible channel for generating these fields involves the pre-conditioning of the circumburst medium: in the early afterglow phase, prompt photons streaming ahead of the GRB external shock can pair produce, seeding the upstream with drifting electron-positron pairs and triggering electromagnetic microinstabilities. To study this process, we employ 2D periodic particle-in-cell simulations in which a cold electron-proton plasma is gradually enriched with warm electron-positron pairs injected at mildly relativistic speeds. We find that continuous pair injection drives the growth of large-scale magnetic fields via filamentation-like instabilities; the temporal evolution of the field is self-similar and depends on a single parameter, $\left[\alpha/(t_{\rm f}\omega_{\rm pi})\right]^{1/2}t\omega_{\rm pi}$, where $\alpha$ is the ratio of final pair beam density to background plasma density, $t_{\rm f}$ is the duration of pair injection, and $\omega_{\rm pi}$ is the plasma frequency of background protons. Extrapolating our results to parameter regimes realistic for long GRBs, we find that upstream pair enrichment generates weak magnetic fields on scales much larger than the proton skin depth; for bright bursts, the extrapolated coherence scale at a shock radius of $R\sim 10^{17}\,{\rm cm}$ is $\lambda_{y}\right>\sim 100\;c/\omega_{\rm pi}$ and the corresponding magnetization is $\sigma\sim 10^{-8}$ for typical circumburst parameters. These results may help explain the persistence of magnetic fields at large distances behind GRB shocks.
@@ 1 Introduction
Our uncertainty regarding the nature of GRB afterglows is intimately tied to our incomplete understanding of relativistic, weakly magnetized, collisionless shocks (e.g., Sironi et al. 2013, 2015; Vanthieghem et al. 2020). The low densities of the environments surrounding GRB progenitors yield Coulomb mean free paths that far exceed the system size, meaning that GRB external shocks are mediated not by binary particle collisions but by collective plasma interactions (Moiseev & Sagdeev, 1963; Medvedev & Loeb, 1999; Gruzinov & Waxman, 1999; Lyubarsky & Eichler, 2006; Brainerd, 2000; Wiersma & Achterberg, 2004; Kato, 2007). The ambient magnetic fields of circumburst environments are typically very weak, with $\sigma_{0}\equiv B_{0}^{2}/4\pi n_{\rm i}m_{\rm i}c^{2}\sim 10^{-9}-10^{-5}$, where $n_{\rm i}$ is the proton density, $m_{\rm i}$ is the proton mass and $c$ is the speed of light. In this weakly magnetized, collisionless regime, we expect shocks to be mediated by magnetic turbulence produced by the filamentation (or Weibel) instability (Weibel, 1959; Fried, 1959), which channels the free energy of an anisotropic flow into magnetic energy, effectively generating magnetic fields from scratch (Achterberg & Wiersma, 2007; Achterberg et al., 2007; Bret et al., 2010; Lemoine & Pelletier, 2010, 2011; Rabinak et al., 2011; Nakar et al., 2011; Shaisultanov et al., 2012). Unfortunately, the evolution of this magnetic turbulence is highly nonlinear – with the magnetic fields backreacting on the same particles that produce them – making Weibel-mediated shocks nearly intractable to study analytically (Kirk et al., 2000; Keshet & Waxman, 2005; Keshet et al., 2009; Bret et al., 2013, 2014; Pelletier et al., 2019; Lemoine et al., 2019c, b, a); only fully kinetic numerical simulations can completely capture these nonlinear interactions (Spitkovsky, 2005; Birdsall & Langdon, 1991; Matsumoto & Omura, 1993), but even these simulations are strained by limited computational resources, leaving many questions still unanswered (Keshet et al., 2009).
@@
A crucial question in the modeling of GRB afterglows concerns the extent of the afterglow emission region. If the magnetic fields powering the synchrotron afterglow are indeed generated by the filamentation instability, then theory predicts that these fields should grow on scales comparable to the proton skin depth (Medvedev & Loeb, 1999), leading to fast decay downstream of the shock (Chang et al., 2008; Gruzinov, 2001; Lemoine, 2015); instead, observations combined with analytical models suggest that these downstream fields may extend up to $10^{8}$ proton skin depths behind the shock (Waxman, 2006). In addition, if particles at the shock were to scatter in exclusively microscale fields, their non-thermal acceleration would be limited (Sironi et al., 2013; Reville & Bell, 2014; Huang et al., 2022). One possible solution to this problem involves the feedback of particles accelerated at the shock (Keshet et al., 2009; Grošelj et al., 2024): as particles are accelerated to higher and higher energies via the Fermi process (Blandford & Eichler, 1987), they will penetrate farther into the upstream, seeding progressively larger-scale magnetic fluctuations that decay more slowly and scatter particles more efficiently. An alternative solution – particularly relevant to long bursts, in which the GRB’s external shock propagates through the dense wind of the Wolf-Rayet progenitor star – involves the loading of the circumburst medium with electron-positron pairs prior to the passage of the shock (Thompson & Madau, 2000; Beloborodov, 2002; Kumar & Panaitescu, 2004; Ramirez-Ruiz et al., 2007; Derishev & Piran, 2016; Grošelj et al., 2022). Early in the afterglow phase (i.e., within a few minutes of the burst), prompt photons stream just ahead of the shock front, where a small fraction could scatter off of electrons in the upstream plasma; the scattered photons are then susceptible to collisions with other outgoing prompt photons, producing warm electron-positron pairs. [^1] The photons required for upstream pair production may alternatively be sourced by synchrotron radiation downstream of the shock (Derishev & Piran, 2016). As each pair increases the optical depth of the upstream medium – thus increasing the probability of photon scattering – the rate of pair creation increases exponentially, ultimately resulting in a pair density profile that decreases away from the shock front (Beloborodov et al., 2014). The streaming of these electron-positron pairs through the circumburst plasma should trigger the filamentation instability well before the shock arrives, potentially pre-seeding the upstream with magnetic fields; if these fields are coherent on sufficiently large scales, they could survive far into the downstream.
@@
In this paper, we overcome the limitations of Garasev & Derishev (2016) by carrying out periodic, 2D, fully kinetic simulations of the continuous enrichment of an electron-proton plasma (with realistic mass ratio) by relativistic electron-positron pairs carrying net momentum in the fluid frame. We also explore the physics of this system in depth, identifying trends that allow us to extrapolate our simulation results to regimes realistic for long GRBs. We conduct an extensive parameter scan over two key quantities: $t_{\rm f}$ – the pair injection duration (or, equivalently, the time for the upstream fluid element to encounter the shock front) – and $\alpha$ – the ratio of the final pair beam density to the background plasma density. In the range of $t_{\rm f}$ and $\alpha$ that we explore, we find that continuous pair injection drives the growth of large-scale magnetic fields via filamentation-like instabilities; the temporal evolution of the system is self-similar and depends on a single parameter, $\left[\alpha/(t_{\rm f}\omega_{\rm pi})\right]^{1/2}t\omega_{\rm pi}$, where $\omega_{\rm pi}$ is the plasma frequency of the background protons. At the end of the pair injection phase, the mean wavenumber of the magnetic field (in units of $\left(c/\omega_{\rm pi}\right)^{-1}$) scales approximately with $\alpha^{-1/12}\,\left(t_{\rm f}\omega_{\rm pi}\right)^{-2/3}$ and the self-generated magnetic energy density (in units of the proton rest mass energy density) scales with $\alpha^{5/8}\,\left(t_{\rm f}\omega_{\rm pi}\right)^{-3/4}$. Using these trends to extrapolate to realistic regimes, we find that upstream pair enrichment generates weak magnetic fields on scales much larger than the proton skin depth; for bright bursts, the extrapolated coherence scale at a shock radius of $R\sim 10^{17}\,{\rm cm}$ is $\lambda_{y}\right>\sim 100\;c/\omega_{\rm pi}$ and the corresponding magnetization is $\sigma\sim 10^{-8}$ for typical circumburst parameters. Since larger-scale fields decay more slowly, our results may help explain the persistence of magnetic fields at large distances behind GRB shocks.
This paper is structured as follows: in Section 2, we detail our simulation setup; in Section 3, we analyze an example simulation, highlighting the salient features of the magnetic field evolution; in Section 4, we compare how the magnetic field evolution varies with $\alpha$ and $t_{\rm f}$; in Section 5, we extrapolate our simulation results to realistic regimes; and in Section 6, we summarize our main findings and conclude. We provide supplementary information in the appendices: in Appendix A, we compare our 2D results with a 3D simulation; in Appendix B, we show results from simulations with reduced mass ratios to confirm that our aforementioned extrapolations are fully justified; in Appendix C, we compare simulations with varying beam Lorentz factors; in Appendix D, we demonstrate that changing our definition of the dominant magnetic wavenumber negligibly changes our conclusions; and in Appendix E, we show that the results presented in this paper are converged numerically.
@@ 2 Simulation Setup
We employ the electromagnetic particle-in-cell (PIC) code TRISTAN-MP (Spitkovsky et al., 2019; Spitkovsky, 2005), which is especially well-suited for handling relativistic flows in collisionless plasmas. The code solves the Maxwell-Vlasov system using a second-order Finite-Difference Time-Domain (FDTD) scheme on a standard Yee mesh (Yee, 1966) and pushes particles using the Boris algorithm (Boris, 1970) and first-order shape functions. The electromagnetic fields are extrapolated to particle positions using a bilinear interpolation function (or a trilinear interpolation function for our 3D run). At each time step, after depositing the electric current to the grid using the charge-conserving zig-zag scheme (Umeda et al., 2003), we apply twenty passes of a 3-point (1-2-1) digital current filter in each direction to smooth out non-physical short-wavelength oscillations.
@@
We run fully periodic 2D simulations (in the $xy$-plane) of a cold, initially unmagnetized electron-proton plasma (the “background”) that is gradually enriched with warm, mildly relativistic electron-positron pairs (the “beam”). The background particles and the beam particles are both distributed homogeneously throughout the simulation box. The background is initially at rest and follows a Maxwellian with $\frac{k_{\rm B}T_{\rm i}}{m_{\rm i}c^{2}}=\frac{m_{\rm e}}{m_{\rm i}}\frac{k_{\rm B}T_{\rm e,\,bg}}{m_{\rm e}c^{2}}=10^{-7}$ (where $k_{\rm B}$ is Boltzmann’s constant, $c$ is the speed of light, $T_{\rm i}$ is the proton temperature, $T_{\rm e,\,bg}$ is the background electron temperature, $m_{\rm i}$ is the proton mass, $m_{\rm e}$ is the electron mass, and $m_{\rm i}/m_{\rm e}=1836$). The beam pairs are drawn from a Maxwell-Jüttner distribution with $\frac{k_{\rm B}T_{\rm b}}{m_{\rm e}c^{2}}=1$ (for $T_{\rm b}$ the initial temperature of both beam species), drifting along the positive $x$ direction with bulk Lorentz factor $\gamma_{\rm b}=1.5$.
@@
The beam particles are injected in pairs at a constant rate throughout the duration of the simulation. The injection rate is determined by two key parameters: $t_{\rm f}$ – the injection duration – and $\alpha$ – the ratio of the total beam density to the total background density at the end of injection. In terms of these quantities, $n_{\rm bg}\times\frac{\alpha}{t_{\rm f}\omega_{\rm pb}}$ electron-positron pairs are injected per plasma time ($\omega_{\rm pb}^{-1}$). Here, $n_{\rm bg}$ is the apparent number density of the background plasma (i.e., the total number density of the background electrons and the protons) and $\omega_{\rm pb}$, the final beam electron plasma frequency, is defined to be
$\omega_{\rm pb}\equiv\sqrt{\frac{4\pi(n_{\rm b}/2)e^{2}}{m_{\rm e}}}\,,$ (1)
@@
for $n_{\rm b}$ the total beam density at the end of pair injection. It is also useful to define the plasma frequency of the background protons,
$\omega_{\rm pi}\equiv\sqrt{\frac{4\pi(n_{\rm bg}/2)e^{2}}{m_{\rm i}}}=\omega_{\rm pb}\sqrt{\frac{m_{\rm e}}{m_{\rm i}\alpha}}\,.$ (2)
@@
In our simulations, all quantities are normalized to the properties of the pair beam at the end of injection to ensure that all length-scales are well-resolved by our grid; since we only consider $\alpha\geq 1$, the skin depth of the protons ($c/\omega_{\rm pi}$) and the skin depth of the background electrons ($c/\omega_{\rm pe}\equiv\sqrt{m_{\rm e}/m_{\rm i}}\,c/\omega_{\rm pi}$) are always at least as large as the skin depth of the beam electrons. Throughout this paper, we cite results both in units of the beam electrons and in units of the background protons.
@@
In our suite of simulations, we consider values of $\alpha$ equal to 1, 2, 4, 8, and 16; for each $\alpha$, we vary $t_{\rm f}$ among $t_{\rm f}\omega_{\rm pb}=5\times 10^{3}$, $10^{4}$, $2\times 10^{4}$, and $4\times 10^{4}$ (respectively corresponding to $t_{\rm f}\omega_{\rm pi}\sim 117/\sqrt{\alpha}$, $233/\sqrt{\alpha}$, $467/\sqrt{\alpha}$, and $934/\sqrt{\alpha}$). These values are small compared to those appropriate for real GRBs, since our realistic mass ratio prevents us from directly simulating larger $\alpha$ and $t_{\rm f}$. However, our range of $\alpha$ and $t_{\rm f}$ is broad enough to allow for reliable extrapolation to higher values.
@@
For our production runs, we primarily use 36 final beam particles per cell per species; the number of background particles per cell per species is thus $36/\alpha$. For a few representative cases, we also use 72 final beam particles per cell per species (and $72/\alpha$ background particles per cell per species) to check for numerical convergence (see Appendix E). We resolve the final beam electron skin depth, $c/\omega_{\rm pb}$, using 10 cells. We employ square simulation boxes with side lengths of either 600 $c/\omega_{\rm pb}$ (6000 cells) or 1000 $c/\omega_{\rm pb}$ ($10^{4}$ cells); we use the larger boxes for our $t_{\rm f}\omega_{\rm pb}=2\times 10^{4}$ and $4\times 10^{4}$ cases to ensure that the dominant scale of the magnetic field is not artificially limited by the box size. For our $\alpha=16$ simulations (i.e., the cases with the largest proton skin depth), $c/\omega_{\rm pi}\sim 200\;c/\omega_{\rm pb}$, so our simulation boxes always fit at least three proton skin depths in each direction.
@@
To check our 2D results, we run a single 3D simulation with our fiducial parameters, $\alpha=2$ and $t_{\rm f}\omega_{\rm pb}=10^{4}$; this simulation is analyzed in Appendix A. For this 3D run, we use 18 final beam particles per cell per species, we resolve $c/\omega_{\rm pb}$ with 10 cells, and we employ a cubic simulation box with a 100 $c/\omega_{\rm pb}$ (1000 cell) side length.
@@
In Appendix E, we show that our simulations converge numerically with respect to both particles per cell and spatial resolution, and we demonstrate that the growth of the magnetic field is not impeded by the size of our simulation box.
3 Case study: $\alpha=2,\;t_{\rm f}\omega_{\rm pb}=10^{4}$